OpenStudy (lgbasallote):

is it possible to turn differential equation general solution into differential equation?

OpenStudy (unklerhaukus):

differentiate ?

OpenStudy (lgbasallote):

just differentiate?

OpenStudy (lgbasallote):

lol @saifoo.khan shut up :p @UnkleRhaukus can you provide an example?

OpenStudy (unklerhaukus):

yep the arbitrary constant will go

OpenStudy (unklerhaukus):

i sure can give me a second

OpenStudy (lgbasallote):

let's say i have \[x^2y + x + c = 0\] dow do i turn that into a differential equation?

OpenStudy (lgbasallote):

let's just use my example :S

OpenStudy (saifoo.khan):

loooooooooool. @lgbasallote

OpenStudy (lgbasallote):

i think the original equation was \[(2xy + 1)dx + (2xy)dy = 0\] or something like that...how do i produce it?

OpenStudy (unklerhaukus):

\[(x^2-y^2)\text dx+(2xy)\text d y=0\] \[\frac{\text d y}{\text dx}=\frac{y^2-x^2}{2xy}\]\[\frac{\text d y}{\text dx}=\frac{\left(\frac yx\right)^2-1}{2\left(\frac yx\right)}\] \[\text{let } \frac y x = v\]\[y=vx\] \[\frac{\text d y }{\text d x}=v+x\frac{\text d v}{\text dx}\]\[v+x\frac{\text d v}{\text dx}=\frac{v^2-1}{2v}\]\[x\frac{\text d v}{\text dx}=\frac{v^2-1}{2v}-v\]\[x\frac{\text d v}{\text dx}=\frac{v^2-1-2v^2}{2v}\]\[x\frac{\text d v}{\text dx}=-\frac{1+v^2}{2v}\]\[\frac{2v}{1+v^2}\text dv=-\frac{\text d x}{ x}\]\[\int\frac{2v}{1+v^2}\text dv=\int-\frac{\text d x}{ x}\]\[\ln(1+v^2)=-\ln x + c\]\[\ln(1+v^2)=\ln e^c-\ln x \]\[\ln(1+v^2)=\ln \frac{k}{x}\]\[1+v^2=\frac kx\]\[1+ \left(\frac yx\right)^2 =\frac kx\] \[x^2+y^2=kx\]

OpenStudy (unklerhaukus):

\[x+\frac {y^2}x=k\] \[\frac{\text d}{\text dx}\left(x+\frac{y^2}{x}\right)=\frac{\text d}{\text dx}\left(k\right)\]\[1-\frac{y^2}{x^2}+2\frac yx\frac{\text dy}{\text dx}=0\]\[1+2\frac yx\frac{\text dy}{\text dx}=\frac{y^2}{x^2}\]\[x^2+2xy\frac{\text dy}{\text dx}=y^2\]\[2xy\frac{\text d y }{\text dx}=y^2-x^2\] \[\frac{\text d y}{\text dx}=\frac{y^2-x^2}{2xy}\]

OpenStudy (lgbasallote):

that...was....TOTALLY WICKED!!!`

OpenStudy (lgbasallote):

awesome...just awesome :O

OpenStudy (lgbasallote):

you used product rule on the derivative of y^2/x right?

OpenStudy (unklerhaukus):

\[x^2y + x + c = 0\] \[y=cx^{-2}-x^{-1}\] \[\frac{\text dy}{\text dx}=\frac{\text d}{\text dx}\left(cx^{-2}-x^{-1}\right)\]

OpenStudy (lgbasallote):

hmm what's the derivative of c/x^2?

OpenStudy (lgbasallote):

-2c/x^3?

OpenStudy (lgbasallote):

but since -2C is still a constant it's just c? so c/x^3?

OpenStudy (lgbasallote):

then -1/x becomes 1/x^2

OpenStudy (unklerhaukus):

thats right i did use the product rule for \(\frac{\text d}{\text dx}\left(\frac{y^2}{x}\right)\)

OpenStudy (lgbasallote):

i dont think this last one is making sense...

OpenStudy (lgbasallote):

ahh sowhat next

OpenStudy (unklerhaukus):

hmm the constant didn't do away like it wasn't ment to

OpenStudy (lgbasallote):

hmm i need to sleep...ill get back to this in the morning -_- lol

OpenStudy (unklerhaukus):

maybe this way is better \[x^2y+x+c=0\]\[x^2y+x=-c\] \[\frac{\text d}{\text dx}\left(x^2y+x\right) =0\]

OpenStudy (unklerhaukus):

using the product rule\[2xy+x^2\frac{\text dy}{\text dx}+1=0\]

OpenStudy (unklerhaukus):

\[x^2\frac{\text dy}{\text dx}=-(1+2xy)\] \[\frac{\text dy}{\text dx}=\frac{-(1+2xy)}{x^2}\] \[{x^2}{\text dy}=-{(1+2xy)}\text dx\] \[{(2xy+1)}\text dx+(x^2){\text dy}=0\]

OpenStudy (unklerhaukus):

thank you

OpenStudy (lgbasallote):

lol nice @UnkleRhaukus i just gained new respect for you <tips hat>